# GRVFL-2V: Graph Random Vector Functional Link Based on Two-View Learning

### v1

@@ Abstract
The classification performance of the random vector functional link (RVFL), a randomized neural network, has been widely acknowledged. However, due to its shallow learning nature, RVFL often fails to consider all the relevant information available in a dataset. Additionally, it overlooks the geometrical properties of the dataset. To address these limitations, a novel graph random vector functional link based on two-view learning (GRVFL-2V) model is proposed. The proposed model is trained on multiple views, incorporating the concept of multiview learning (MVL), and it also incorporates the geometrical properties of all the views using the graph embedding (GE) framework. The fusion of RVFL networks, MVL, and GE framework enables our proposed model to achieve the following: i) efficient learning: by leveraging the topology of RVFL, our proposed model can efficiently capture nonlinear relationships within the multi-view data, facilitating efficient and accurate predictions; ii) comprehensive representation: fusing information from diverse perspectives enhance the proposed model’s ability to capture complex patterns and relationships within the data, thereby improving the model’s overall generalization performance; and iii) structural awareness: by employing the GE framework, our proposed model leverages the original data distribution of the dataset by naturally exploiting both intrinsic and penalty subspace learning criteria. The evaluation of the proposed GRVFL-2V model on various datasets, including 27 UCI and KEEL datasets, 50 datasets from Corel5k, and 45 datasets from AwA, demonstrates its superior performance compared to baseline models. These results highlight the enhanced generalization capabilities of the proposed GRVFL-2V model across a diverse range of datasets.
@@ 1 Introduction
The artificial neural networks (ANNs) belong to the class of non-parametric learning methods that are used for estimating or approximating functions that may depend on a large number of inputs and outputs [1]. Inspired by the topology of biological neural networks, ANNs have found applications in diverse domains, such as image recognition [2], Alzheimer’s disease diagnosis [3], stock price prediction [4], brain age prediction [5], and so on. The efficacy of ANNs depends upon several factors, including the quality and quantity of data, computational resources, and the effectiveness of underlying algorithms [6]. The ANNs are trained by optimizing a cost function, which quantifies the disparity between model predictions and actual observations. The adjustment of the network’s weights and biases is facilitated through back-propagation, employing an iterative technique known as the gradient descent (GD) method. To enhance the agreement between predicted outcomes and actual observations, the optimization process requires fine-tuning the network’s weights and biases. Nevertheless, GD-based algorithms come with inherent limitations, including slow convergence [7], difficulty in achieving global minima [8] and heightened sensitivity to the selection of learning rate and the point of initialization.
@@
Randomized neural networks (RNNs) [9] have emerged as a unique solution to overcome the difficulties faced in training ANNs using gradient descent (GD) methods. Unlike conventional ANNs, RNNs adopt a distinct approach where the network’s weights and biases are randomly selected from a specified range and remain constant throughout the training process. The determination of output layer parameters in RNNs is achieved through a closed-form solution [10], as opposed to iterative-based optimization techniques commonly employed in traditional ANNs. The prominent examples of RNNs include the random vector functional link (RVFL) network [11, 12] and the extreme learning machines (ELM) [13]. RVFL has garnered significant attention within the realm of RNNs due to its distinctive features, including direct connections between input and output layers that improve generalization performance, a straightforward architecture aligned with Occam’s Razor principle and PAC learning theory [14], a reduced number of parameters, and the universal approximation capability [15]. Furthermore, the direct connections help to regulate the randomness in RVFL [1, 16], while the output parameters are computed analytically through methods like pseudo-inverse or least-square techniques.
@@
Various enhancements have been implemented to optimize the learning capabilities of RVFL, with notable advancements such as the sparse pre-trained random vector functional link (SP-RVFL) network [17] and the discriminative manifold RVFL neural network (DM-RVFLNN) [18]. The SP-RVFL network utilizes diverse weight initialization methods to improve the generalization performance of standard RVFL models, while the DM-RVFLNN integrates manifold learning with a soft label matrix to enhance the model’s discriminative capacity by increasing the margin between samples from different classes. In addition, techniques such as manifold learning based on an in-class similarity graph have been employed to enhance the compactness and similarity of samples within the same class. To address noise and outliers, Cui et al. [19] proposed an RVFL-based approach that incorporates a novel feature selection (FS) technique, enhancing the efficiency and robustness of RVFL through the augmented Lagrangian method. Furthermore, the integration of a kernel function into RVFL has led to the development of the kernel exponentially extended RVFL network (KERVFLN) [20], demonstrating the adaptability and potential of RVFL in various machine learning applications. Recent advancements in RVFL theory include RVFL+, kernel RVFL+ (KRVFL+) [21], incremental learning paradigm with privileged information for RVFL (IRVFL+), and improved fuzziness-based RVFL [22, 23], showcasing the ongoing evolution and versatility of RVFL models in addressing diverse machine-learning challenges, including multilabel classification tasks using RNNs. Some recent advancements in RVFL theory include neuro-fuzzy RVFL for classification and regression problems [16], kernel ridge regression-based randomized network for brain age classification and estimation [24], domain-incremental learning without forgetting based on RVFL [25], etc.
@@
The fusion of the graph embedding (GE) [30] framework with the RVFL model has played a crucial role in enhancing the learning process of RVFL. By incorporating the GE framework, the RVFL model is able to capture the geometric relationships present within the dataset, which were previously overlooked [12]. To incorporate the geometrical relationships within a dataset, Malik et al. [31] introduced the graph-embedded intuitionistic fuzzy weighted RVFL (GE-IFWRVFL) model. This model not only integrates the intuitionistic fuzzy (IF) membership scheme to handle noisy data and outliers but also preserves the geometric characteristics through the GE framework. The fusion of IF and GE with RVFL has proven to enhance the resilience and performance of the RVFL significantly. Recent advancements in RVFL have seen the utilization of GE framework in extended graph-embedded RVFL (EGERVFL) and graph-embedded intuitionistic fuzzy RVFL network for class imbalance learning (GE-IFRVFL-CIL) as proposed in [32] and [33] respectively. The experimental results from [31, 32, 33], along with studies in [12], strongly support the notion that the performance of RVFL is greatly improved through the fusion of the GE framework. This emphasizes the importance of considering the geometric attributes of a dataset in the learning process to enhance the overall performance of RVFL.
@@
RVFL is categorized as a shallow learning algorithm due to its single hidden layer, which may potentially hinder its ability to fully grasp complex patterns and subtle nuances in a dataset. While RVFL may excel in specific classification tasks, its architecture could limit its capability to accurately interpret intricate patterns or extract detailed features from the dataset. The simplicity of RVFL’s architecture offers advantages in terms of computational efficiency and ease of implementation. However, when confronted with datasets containing intricate patterns or nuanced relationships among variables, more sophisticated learning frameworks are necessary to effectively capture the underlying complexities inherent in the dataset. To mitigate the limitations imposed by the subsurface learning nature of RVFL, presenting the same information from multiple perspectives can offer a solution. By training RVFL on multiple views of the data, it becomes possible to compensate for its inherent learning structure and enhance its ability to comprehend intricate patterns and subtle nuances within the dataset. Hence, leveraging the multiview learning (MVL) framework can significantly enhance the learning efficiency of RVFL.
@@
Recognizing the importance of a dataset’s inherent geometrical characteristics and MVL, we propose the novel graph random vector functional link based on two-view learning (GRVFL-2V). The GRVFL-2V method incorporates the intrinsic and penalty graphical representations of multiview datasets within the GE framework, thereby capturing the geometrical properties of the multiview data. By fusing the MVL and GE framework with RVFL, the learning capabilities of RVFL are greatly enhanced, allowing it to effectively tackle classification challenges. In order to ensure simplicity and effectiveness, the proposed GRVFL-2V model incorporates two views at a time. This strategy allows for a balance between complexity and performance, as it utilizes the complementary information from these two views to enhance effectiveness while still maintaining simplicity. By leveraging information from multiple views, GRVFL-2V significantly improves the classification performance of RVFL by integrating both geometric and discriminative information from both views using intrinsic and penalty-based subspace learning criteria within the GE framework.
@@
    We present a generic framework that fuses random vector functional link (RVFL) with multiview learning (MVL) and graph embedding (GE) framework. This novel model is called the graph random vector functional neural network based on two-view learning (GRVFL-2V).
@@
    The proposed GRVFL-2V model is developed upon the foundation of the RVFL architecture, which is known for its simplicity and efficiency. However, it goes beyond the limitations of shallow learning of RVFL by integrating the concept of multiview learning (MVL). By leveraging multiple views, the model aims to enhance classification performance using multiple perspectives of the dataset.
@@
    The proposed GRVFL-2V model integrates geometrical information from multiview data by embedding intrinsic and penalty subspace learning (SL) criteria within the GE framework. It employs local Fisher discriminant analysis (LFDA) [34] and graph regularization parameters to effectively utilize the GE framework.
@@
    Our experiments encompassed 27 UCI and KEEL datasets, 50 datasets from Corel5k, and 45 datasets from AwA. Through comprehensive statistical analyses, we demonstrate the superior generalization performance of our proposed model when compared to baseline models.
The paper is organized as follows: Section 2 provides a comprehensive overview of preliminary works. The mathematical derivation of the proposed model is presented in Section 3, while Section 4 discusses the computational complexity of the proposed model. Section 5 presents the experimental results and statistical analyses. Finally, Section 6 concludes the paper and provides recommendations for future research.
@@ 2.1 Notations
Let the two-view training dataset $S=\bigl{\{}(x_{i}^{A},x_{i}^{B},y_{i})|x_{i}^{A}\in\mathbb{R}^{1\times n},x_{i}^{B}\in\mathbb{R}^{1\times m},\,y_{i}\in\{-1,1\},i=1,2,\cdots,l\bigl{\}}$, where $x_{i}^{A}$ and $x_{i}^{B}$ represents the feature vector of each data sample corresponding to $view-A$ and $view-B$, respectively and $y_{i}$ represents the label of the individual data for both the views. Let $X_{A}\in\mathbb{R}^{l\times n}$ and $X_{B}\in\mathbb{R}^{l\times m}$ be the input matrices for $view-A$ and $view-B$, respectively and $Y_{true}\in\mathbb{R}^{l\times 2}$ represents the one hot encoding matrix of the output. $(\cdot)^{t}$ represents the transpose operator.
@@ 2.4 Literature Review on Multiview Learning (MVL)
Multiview learning (MVL) is an emerging area of research that holds significant promise in enhancing the generalization performance across various learning tasks. By integrating multiple feature sets, each offering unique and complementary insights, MVL has the potential to significantly improve overall model performance [36, 37, 38]. The abundance of diverse data types in practical applications has led to the development of MVL. In real-world scenarios, samples from different perspectives may reside in distinct spaces or exhibit vastly different distributions, often due to significant disparities between views [39, 40]. However, conventional approaches typically address such data by employing a cascade strategy, wherein multiview data is amalgamated into a single-view format through the concatenation of heterogeneous feature spaces into a homogeneous one. However, this cascading approach overlooks the distinctive statistical properties of each view and is plagued by the curse of dimensionality problems. One notable advantage of MVL is that it can enhance the performance of a standard single-view approach by leveraging manually generated multiple views.
@@
In MVL, a distinct function is developed for each view, with the overall goal of constructing a unified function to optimize all individual functions jointly. This approach enhances the generalization performance of the framework across multiple views. According to Zhao et al. [36], MVL models can be categorized into three main groups: co-training style algorithms, co-regularization style algorithms, and margin consistency style algorithms. Co-training style algorithms focus on enhancing mutual agreement among diverse views. Conversely, co-regularization style algorithms aim to minimize discrepancies during the learning process. MVL learning has been successfully implemented in various hyper-plane based classifiers, such as SVM-2K [41], multiview twin support vector machine [42] (MvTSVM), deep multi-view multiclass TSVM [43], MVL with robust double-sided TSVM [44], etc. A notable application of MVL has been found in predicting DNA-binding protein using RVFL [45]. Furthermore, MVL has demonstrated its effectiveness across various application scenarios, including enhancing image classification, annotation, and retrieval performance [46], predicting financial distress [47], forecasting multiple stages of Alzheimer’s Disease progression [48], and identifying product adoption intentions from social media data [49]. MVL is now widely used in many various domains [50, 51] and research endeavors.
@@
3 The Proposed Graph Random Vector Functional Link Neural Network based on Two-View Learning (GRVFL-2V)
@@
In this section, we introduce a novel GRVFL-2V, which fuses the RVFL with MVL. Our proposed model is structured to leverage the strengths of RVFL while incorporating both intrinsic and penalty graphical representations of multiview data through the GE framework. To derive the network output weights, the optimization process integrates subspace learning (SL) criteria, incorporating both intrinsic and penalty-based SL methodologies within the GE framework. To preserve the geometric structure effectively, we incorporate local Fisher discriminant analysis (LFDA) [34] into the GE framework. Additionally, we introduce a regularization parameter for GE to enhance the learning process further. To trade off the error between multiple views, we introduce a coupling term in the primal optimization problem of the proposed model. This term minimizes the combined error from both views, resulting in an improved generalization performance for our model. The proposed optimization problem of the GRVFL-2V model is presented below:
@@
where $W_{A}$ and $W_{B}$ are the randomly generated weights matrices, and $b_{A}$ and $b_{B}$ represents the bias column vectors of the bias matrices $B_{A}$ and $B_{B}$, respectively. Since, all the columns of the bias matrix $B_{A}$ are identical, hence $b_{A}$ can be chosen to be any column of $B_{A}$. A similar argument follows for $b_{B}$. The algorithm for the proposed GRVFL-2V is presented in Algorithm 1.

### v3

# GRVFL-MV: Graph Random Vector Functional Link Based on Multi-View Learning

@@ Abstract
The classification performance of the random vector functional link (RVFL), a randomized neural network, has been widely acknowledged. However, due to its shallow learning nature, RVFL often fails to consider all the relevant information available in a dataset. Additionally, it overlooks the geometrical properties of the dataset. To address these limitations, a novel graph random vector functional link based on multi-view learning (GRVFL-MV) model is proposed. The proposed model is trained on multiple views, incorporating the concept of multiview learning (MVL), and it also incorporates the geometrical properties of all the views using the graph embedding (GE) framework. The fusion of RVFL networks, MVL, and GE framework enables our proposed model to achieve the following: i) efficient learning: by leveraging the topology of RVFL, our proposed model can efficiently capture nonlinear relationships within the multi-view data, facilitating efficient and accurate predictions; ii) comprehensive representation: fusing information from diverse perspectives enhance the proposed model’s ability to capture complex patterns and relationships within the data, thereby improving the model’s overall generalization performance; and iii) structural awareness: by employing the GE framework, our proposed model leverages the original data distribution of the dataset by naturally exploiting both intrinsic and penalty subspace learning criteria. The evaluation of the proposed GRVFL-MV model on various datasets, including 29 UCI and KEEL datasets, 50 datasets from Corel5k, and 45 datasets from AwA, demonstrates its superior performance compared to baseline models. These results highlight the enhanced generalization capabilities of the proposed GRVFL-MV model across a diverse range of datasets.
@@ 1 Introduction
The artificial neural networks (ANNs) belong to the class of non-parametric learning methods that are used for estimating or approximating functions that may depend on a large number of inputs and outputs [45]. Inspired by the topology of biological neural networks, ANNs have found applications in diverse domains, such as image recognition [31], Alzheimer’s disease diagnosis [37], stock price prediction [11], and so on. The efficacy of ANNs depends upon several factors, including the quality and quantity of data, computational resources, and the effectiveness of underlying algorithms [1]. The ANNs are trained by optimizing a cost function, which quantifies the disparity between model predictions and actual observations. The adjustment of the network’s weights and biases is facilitated through back-propagation, employing an iterative technique known as the gradient descent (GD) method. To enhance the agreement between predicted outcomes and actual observations, the optimization process requires fine-tuning the network’s weights and biases. Nevertheless, GD-based algorithms come with inherent limitations, including slow convergence [17], difficulty in achieving global minima [10] and heightened sensitivity to the selection of learning rate and the point of initialization.
@@
Randomized neural networks (RNNs) [30] have emerged as a unique solution to overcome the difficulties faced in training ANNs using gradient descent (GD) methods. Unlike conventional ANNs, RNNs adopt a distinct approach where the network’s weights and biases are randomly selected from a specified range and remain constant throughout the training process. The determination of output layer parameters in RNNs is achieved through a closed-form solution [33], as opposed to iterative-based optimization techniques commonly employed in traditional ANNs. The prominent examples of RNNs include the random vector functional link (RVFL) network [24] and the RVFL without direct link (RVFLwoDL) (also known as extreme learning machine (ELM))[14]. RVFL has garnered significant attention within the realm of RNNs due to its distinctive features, including direct connections between input and output layers that improve generalization performance, a straightforward architecture aligned with Occam’s Razor principle and PAC learning theory [18], a reduced number of parameters, and the universal approximation capability [15]. Furthermore, the direct connections help to regulate the randomness in RVFL [45], while the output parameters are computed analytically through methods like pseudo-inverse or least-square techniques.
@@
Various enhancements have been implemented to optimize the learning capabilities of RVFL, with notable advancements such as the sparse pre-trained random vector functional link (SP-RVFL) network [47] and the discriminative manifold RVFL neural network (DM-RVFLNN) [20]. The SP-RVFL network utilizes diverse weight initialization methods to improve the generalization performance of standard RVFL models, while the DM-RVFLNN integrates manifold learning with a soft label matrix to enhance the model’s discriminative capacity by increasing the margin between samples from different classes. In addition, techniques such as manifold learning based on an in-class similarity graph have been employed to enhance the compactness and similarity of samples within the same class. To address noise and outliers, authors in [4] proposed an RVFL-based approach that incorporates a novel feature selection (FS) technique, enhancing the efficiency and robustness of RVFL through the augmented Lagrangian method. Furthermore, similar to support vector machines and their variants [27], incorporating a kernel function into RVFL has resulted in the kernel exponentially extended RVFL network (KERVFLN) [2], demonstrating the adaptability and potential of RVFL in various machine learning applications. Recent advancements in RVFL theory include RVFL+, kernel RVFL+ (KRVFL+) [46], and incremental learning paradigm with privileged information for RVFL (IRVFL+), showcasing the ongoing evolution and versatility of RVFL models in addressing diverse machine-learning challenges, including multilabel classification tasks using RNNs. Some recent advancements in RVFL theory include ensemble deep variant of RVFL based on fuzzy inference system [29], kernel ridge regression-based randomized network for brain age classification and estimation [25], domain-incremental learning without forgetting based on RVFL [21], etc.
@@
The incorporation of the graph embedding (GE) framework [43] into the RVFL model enhances its learning capabilities by embedding the intrinsic geometric structure of the data into the feature space. This integration enables the RVFL model to preserve local and global geometrical relationships, which were previously neglected, thereby improving its representational power and generalization performance.To incorporate the geometrical relationships within a dataset, authors in [22] introduced the graph-embedded intuitionistic fuzzy weighted RVFL (GE-IFWRVFL) model. This model not only integrates the intuitionistic fuzzy (IF) membership scheme to handle noisy data and outliers but also preserves the geometric characteristics through the GE framework. The fusion of IF and GE with RVFL has proven to enhance the resilience and performance of the RVFL significantly. The experimental results from [9], along with studies in [23], strongly support the notion that the performance of RVFL is greatly improved through the fusion of the GE framework. This emphasizes the importance of considering the geometric attributes of a dataset in the learning process to enhance the overall performance of RVFL.
@@
RVFL is categorized as a shallow learning algorithm due to its single hidden layer, which may potentially hinder its ability to fully grasp complex patterns and subtle nuances in a dataset. While RVFL may excel in specific classification tasks, its architecture could limit its capability to accurately interpret intricate patterns or extract detailed features from the dataset. The simplicity of RVFL’s architecture offers advantages in terms of computational efficiency and ease of implementation. However, when confronted with datasets containing intricate patterns or nuanced relationships among variables, more sophisticated learning frameworks are necessary to effectively capture the underlying complexities inherent in the dataset. To address the constraints associated with RVFL’s shallow architecture, adopting a multiview learning (MVL) framework presents a viable solution. By training RVFL on multiple views of the data, the model can integrate diverse perspectives, thereby mitigating its structural limitations and enhancing its ability to discern intricate patterns. Leveraging MVL not only enriches the representational capacity of RVFL but also significantly improves its learning efficiency, making it more suitable for complex data analysis tasks.
@@
Recognizing the importance of a dataset’s inherent geometrical characteristics and MVL, we propose the novel graph random vector functional link based on multi-view learning (GRVFL-MV). The GRVFL-MV method incorporates the intrinsic and penalty graphical representations of multiview datasets within the GE framework, thereby capturing the geometrical properties of the multiview data. By fusing the MVL and GE framework with RVFL, the learning capabilities of RVFL are greatly enhanced, allowing it to effectively tackle classification challenges. In order to ensure simplicity and effectiveness, the proposed GRVFL-MV model incorporates two views at a time. This strategy allows for a balance between complexity and performance, as it utilizes the complementary information from these two views to enhance effectiveness while still maintaining simplicity. By leveraging information from multiple views, GRVFL-MV significantly improves the classification performance of RVFL by integrating both geometric and discriminative information from both views using intrinsic and penalty-based subspace learning criteria within the GE framework.
@@
    We present a generic framework that fuses random vector functional link (RVFL) with multiview learning (MVL) and graph embedding (GE) framework. This novel model is called the graph random vector functional neural network based on multi-view learning (GRVFL-MV).
@@
    The proposed GRVFL-MV model is developed upon the foundation of the RVFL architecture, which is known for its simplicity and efficiency. However, it goes beyond the limitations of shallow learning of RVFL by integrating the concept of multiview learning (MVL). By leveraging multiple views, the model aims to enhance classification performance using multiple perspectives of the dataset.
@@
    The proposed GRVFL-MV model integrates geometrical information from multiview data by embedding intrinsic and penalty subspace learning (SL) criteria within the GE framework. It employs local Fisher discriminant analysis (LFDA) [34] and graph regularization parameters to effectively utilize the GE framework.
@@
    Our experiments encompassed 29 UCI and KEEL datasets, 50 datasets from Corel5k, and 45 datasets from AwA. Through comprehensive statistical analyses, we demonstrate the superior generalization performance of our proposed model when compared to baseline models.
The structure of the paper is as follows: Section 2 offers a detailed overview of the preliminary works that form the foundation of this study. Section 3 delves into the mathematical formulation and computational complexity of the proposed model. Section 4 discusses the experimental results alongside statistical analyses. Section 5 provides an in-depth evaluation of the proposed model’s performance across different dataset types. Finally, Section 6 concludes the paper with key findings and suggestions for future research directions.
@@ 2.1 Notations
For simplicity, we consider the notations w.r.t. two-views. Let the two-view training dataset $S=\bigl{\{}(x_{i}^{A},x_{i}^{B},y_{i})|x_{i}^{A}\in\mathbb{R}^{1\times n},x_{i}^{B}\in\mathbb{R}^{1\times m},\,y_{i}\in\{-1,1\},i=1,2,\cdots,l\bigl{\}}$, where $x_{i}^{A}$ and $x_{i}^{B}$ represents the feature vector of each data sample corresponding to $view-A$ and $view-B$, respectively and $y_{i}$ represents the label of the individual data for both the views. Let $X_{A}\in\mathbb{R}^{l\times n}$ and $X_{B}\in\mathbb{R}^{l\times m}$ be the input matrices for $view-A$ and $view-B$, respectively and $Y_{true}\in\mathbb{R}^{l\times 2}$ represents the one hot encoding matrix of the output. $(\cdot)^{t}$ represents the transpose operator.
@@ 2.4 Literature Review on Multiview Learning (MVL)
Multiview learning (MVL) is an emerging research paradigm with substantial potential to enhance generalization performance across diverse learning tasks. By integrating multiple feature representations, each capturing distinct and complementary information, MVL effectively enriches the learned feature space, leading to improved model robustness and predictive accuracy [36]. The abundance of diverse data types in practical applications has led to the development of MVL. In real-world scenarios, samples from different perspectives may reside in distinct spaces or exhibit vastly different distributions, often due to significant disparities between views [19]. However, conventional approaches typically address such data by employing a cascade strategy, wherein multiview data is amalgamated into a single-view format through the concatenation of heterogeneous feature spaces into a homogeneous one. However, this cascading approach overlooks the distinctive statistical properties of each view and is plagued by the curse of dimensionality problems. One notable advantage of MVL is that it can enhance the performance of a standard single-view approach by leveraging manually generated multiple views.
@@
In MVL, a distinct function is developed for each view, with the overall goal of constructing a unified function to optimize all individual functions jointly. This approach enhances the generalization performance of the framework across multiple views. According to [49], MVL models can be categorized into three main groups: co-training style algorithms, co-regularization style algorithms, and margin consistency style algorithms. Co-training style algorithms focus on enhancing mutual agreement among diverse views. Conversely, co-regularization style algorithms aim to minimize discrepancies during the learning process. MVL learning has been successfully implemented in various hyper-plane based classifiers, such as SVM-2K [8], multiview twin support vector machine [39] (MvTSVM), deep multi-view multiclass TSVM [38], etc. A notable application of MVL has been found in predicting DNA-binding protein using RVFL [28]. Furthermore, MVL has demonstrated its effectiveness across various application scenarios, including enhancing image classification, annotation, and retrieval performance [3], predicting financial distress [35], and identifying product adoption intentions from social media data [48]. MVL is now widely used in many various domains [44] and research endeavors.
@@
3 The Proposed Graph Random Vector Functional Link Neural Network based on Multi-View Learning (GRVFL-MV)
@@
In this section, we introduce a novel GRVFL-MV, which fuses the RVFL with MVL. Our proposed model is structured to leverage the strengths of RVFL while incorporating both intrinsic and penalty graphical representations of multiview data through the GE framework. To derive the network output weights, the optimization process integrates subspace learning (SL) criteria, incorporating both intrinsic and penalty-based SL methodologies within the GE framework. To preserve the geometric structure effectively, we incorporate local Fisher discriminant analysis (LFDA) [34] into the GE framework. Additionally, we introduce a regularization parameter for GE to enhance the learning process further. To trade off the error between multiple views, we introduce a coupling term in the primal optimization problem of the proposed model. This term minimizes the combined error from both views, resulting in an improved generalization performance for our model. The proposed optimization problem of the GRVFL-MV model is presented below:
@@
where $W_{A}$ and $W_{B}$ are the randomly generated weights matrices, and $b_{A}$ and $b_{B}$ represents the bias column vectors of the bias matrices $B_{A}$ and $B_{B}$, respectively. Since, all the columns of the bias matrix $B_{A}$ are identical, hence $b_{A}$ can be chosen to be any column of $B_{A}$. A similar argument follows for $b_{B}$. The algorithm for the proposed GRVFL-MV is presented in Algorithm 1.